OpenStudy (anonymous):

What is the range of the function defined by the equation y = x2 + 4x?

OpenStudy (anonymous):

x for all real numbers

OpenStudy (anonymous):

What does that mean?

OpenStudy (anonymous):

y=x2 + 4x = (x+2)^2 - 4 therefore range of y is [-4,+infinity]

OpenStudy (amistre64):

range tends to mean a span of things; in functions, the range is the span of outcomes possible

OpenStudy (anonymous):

I think I understand, thanks so much! I appreciate it!

OpenStudy (anonymous):

y=x2 + 4x = (x+2)^2 - 4 the minimum value of (x+2)^2 for any real value of x is 0 the maximum value of (x+2)^2 for any real value of x is +infinity therefore range of y is [-4,+infinity]